(5 points) The Highway Safety Department wants to study the driving habits of individuals. A sample of 38 cars traveling on a particular stretch of highway revealed an average speed of 68.9 miles per hour with a standard deviation of 5.9 miles per hour. Round to 4 decimal places.

1.Calculate a 99% confidence interval for the true mean speed of all cars on this particular stretch of highway.

( , )

2. Based on the confidence interval do you think the true average speed is above the speed limit of 65 mph?

because

3. What sample size is needed to estimate the true average speed to within 2 mph at 99% confidence? Note: For consistency's sake, round your t* value to 3 decimal places before calculating the necessary sample size.

Choose n =
Answer #1

1) At 99% confidence interval the critical value is t* = 2.715

The 99% confidence interval for the population mean is

\bar x +/- t* * s/\sqrt n

= 68.9 +/- 2.715 * 5.9/\sqrt 38

= 68.9 +/- 2.5985

= 66.3015, 71.4985

2) Yes the true average speed is above the speed limit of 65 mph, because the true average speed limit lies within the confidence bound 66.3015 and 71.4985.

3) Margin of error = 2

or,  t* * s/\sqrt n = 2

or, 2.715 * 5.9/\sqrt n = 2

or, n = (2.715 * 5.9/2)^2

or, n = 65